OpenStudy (faman39):

Keisha's parents want to save $20,000 in her college savings account over the 15 next years . They have $8,000 to use as an initial deposit. What simple annual interest rate do they need to meet their goal? Round your answer to the nearest tenth. 2.5% 2.7% 10% 12%

OpenStudy (faman39):

@mark_o. can you able to help me with this?

OpenStudy (faman39):

@sam can you able to help with it?

OpenStudy (anonymous):

\[20000=8000(1+x)^{15}\]

OpenStudy (anonymous):

Can you solve for x?

OpenStudy (faman39):

I could not get this :(

OpenStudy (faman39):

2.5 still getting wrong answer

OpenStudy (anonymous):

\[5/2 = (1+x)^{15}\] \[log(5/2)=15log(1+x)\] \[\frac{log(5/2)}{15}=log(1+x)\] \[10^{\frac{log(5/2)}{15}}=1+x\] \[1.062999 =1+x\] x=0.0629 or 6.29% nearest 10th would be 10%

OpenStudy (faman39):

Thank you so much @freewilly922 !

OpenStudy (anonymous):

r=rate 8000+r*15*8000=20,000 r=(20,000-8.000)/(15*8,000) r= 0.10=10%

OpenStudy (faman39):

Thank you so much @mark_o. ! :)

OpenStudy (anonymous):

YW, have fun...:D

OpenStudy (faman39):

Thank you for coming ! :)

OpenStudy (anonymous):

ok, yw..:D